# Deep Submodular Peripteral Networks

### v2

@@ 5.2 Application to Experimental Design
In the context of machine learning, experimental design involves selecting a subset of unlabeled examples from a large pool for labeling to create a training set (Pronzato and Pázman, 2013). In this section, we evaluate the performance of a learnt DSPN based on how well a simple linear model generalizes when trained on a training set that was chosen by a DSPN. We consider an offline setting where assume the full unlabeled pool is available and an online setting where the unlabeled pool is presented to the model in a non-iid stream. To test the robustness of our framework to real-world challenges, we add duplicates such that the ground set is heavily class-imbalanced (the procedure to generate the class-imbalanced set is shown in Appendix H).
@@
Offline Setting: The offline setting assumes that the full unlabeled pool is accessible by the summarization procedure. The results of these experiments are presented in Figure 6. Our approach, denoted as DSPN in the figure, applies greedy maximization to the learnt DSPN to generate the summary. We also consider a cheating experiment by applying greedy to the target FL after removing duplicates from the ground set; we refer to this approach as Target FL. Finally, we consider three baseline summarization procedures, CRAIG (Mirzasoleiman et al., 2020), K-centers (Sener and Savarese, 2017), and random subset selection as done in Okanovic et al. (2024). We find that the learnt DSPN achieves performance comparable to the target FL in all datasets up to a budget of 100. Beyond that, the difference between DSPN and the Target FL grows but the DSPN continues to generate higher quality than any of the baselines.
@@
Online Setting: We consider an online setting where the data is presented to the model in a non-iid, class-incremental stream, as done in the continual learning literature (Lopez-Paz and Ranzato, 2017; Chaudhry et al., 2019; Aljundi et al., 2019; Zhou et al., 2023). Unlike the facility location function which requires the full ground set for evaluation, the DSPN can be maximized in a streaming manner. Thus, we employ the streaming maximization algorithm proposed by Lavania et al. (2020) to maximize the DSPN, due to its effectiveness in real-world settings and its lack of additional memory requirements. This approach is referred to as DSPN-online in Figure 6.
@@
We compare our approach against two streaming summarization baselines. The first approach, reservoir sampling (Vitter, 1985), is an algorithm that enables random sampling of a fixed number of items from a data stream as if the sample was drawn from the entire population at once. The second approach, VeSSAL (Saran et al., 2023), seeks to select samples such that the summary maximizes the determinant of the gram matrix constructed from the model embeddings. We also apply (offline) greedy maximization to the target facility location (Target FL-offline) and the learned DSPN (DSPN-offline), which serve as upper bounds to what is attainable by streaming DSPN maximization.
@@
In Figure 6, we present quantitative results where a linear model is trained on a given streaming summary after labels have been obtained. We find that DSPN-online significantly outperforms the other streaming baselines on all datasets. Surprisingly, DSPN-offline and DSPN-online achieve similar performance, demonstrating the flexibility of DSPN. FL-offline tends to generate the highest quality summaries, but requires access to the full ground set and assumes duplicates can be manually removed.
@@ 5.3 Ablations
Table 1: Ablations on Imagenet100. The importance of including feedback, using learnt features, and enforcing submodularity. “w/o feedback" indicates a DSPN model trained without feedback, while “w/o target feedback” is when the oracle is only queried. For Set Transformer, streaming maximization fails to produce a summary of the required size, therefore, we omit that result. All results are reported at a fixed set size of 100.
In Table 1, we provide Imagenet100 results on sets of size 100 that explore the importance of various components of the DSPN. We report three numbers for each ablation: (1) FL Eval which corresponds to the normalized FL evaluation discussed in § 5.1 (2) Offline Acc which denotes the final accuracy of a linear model in the offline experimental design setting and (3) Online Acc denotes the final accuracy of a linear model in the online experimental design setting.
Removing (FL) Feedback: Active selection is obtained by maximizing the DSPN or Target FL during training (denoted as DSPN feedback and Target feedback respectively). We find that only removing the sets obtained from target feedback has little to no impact on the final performance of the DSPN, suggesting that it is not necessary to optimize the target function. This shows that our framework can be applied to learn DSPNs even when it is not easy to optimize the target. However, we find that completely removing Active selection significantly hurts the final DSPN.
Removing Learnable Pillar: We assess the utility of a learnable pillar, by training a DSPN with a frozen pillar. In this case, we attempt to learn the weights of a DSF on random projections of CLIP features. We discover that the DSPN without learnable features significantly underperforms its counterpart. We discuss this further in appendix I.
Removing Submodularity: We study the utility of submodular constraints to learn the target. To this end, we compare DSPN against a Deep Set (Zaheer et al., 2017) and SetTransformer (Lee et al., 2019) architecture in the third section of table 1. Upon doing so, we discover that the absence of submodularity hampers the performance for all the metrics that we consider, since the greedy algorithm has no guarantees for an arbitrary set function.
@@ 6 Conclusion and Future Work
In this paper, we propose DSPNs and methods for training them. Since the computational cost of querying a DSPN is independent of dataset size $n$, and the training scheme leverages automatically generated $(E,M)$ pairs, our method learns practical submodular functions in a scalable manner. In the future, it may be possible to scale our training framework to massive heterogeneous datasets to develop foundation DSPNs that have strong zero-shot summarization capabilities on a wide array of domains.
We also propose the peripteral loss which leverages numerically graded relationships between pairs of objects. In our framework, the graded pairwise preferences are provided from an oracle function and are used to learn a submodular function that captures these relationships. However, the scope of the peripteral loss extends far beyond training DSPNs and has many applications in modern machine learning such as learning a reward model in the context of RLHF.
7 Acknowledgements
We thank Suraj Kothawade, Krishnateja Killamsetty and Rishabh Iyer for several earlier discussions. This work is supported in part by the CONIX Research Center, one of six centers in JUMP, a Semiconductor Research Corporation (SRC) program sponsored by DARPA, NSF Grant Nos. IIS-2106937 and IIS-2148367, and by NIH/NHGRI U01 award HG009395.
@@ Appendix A Other Related Work
Learning Non-submodular Set functions
@@
Recently, there has been a growing focus on leveraging data-driven neural methods for estimating set functions, with practical applications spanning anomaly detection, amortized clustering of mixture models, and point cloud classification [Qi et al., 2017]. Common to these applications is the fundamental approach of embedding each element within a set into dense embeddings, followed by permutation-invariant aggregation (set-to-vector transformation), and subsequent integration into a deep neural network [Zaheer et al., 2017, Horn et al., 2020, Ravanbakhsh et al., 2017]. The significance of incorporating interaction among set elements has been emphasized by Lee et al. [2019], who argue that conventional aggregations, such as average-pooling, can be interpreted as a specialized form of attention. Consequently, they propose an attention-based architecture known as the set-transformer. Building upon this, Mialon et al. [2020] have made computational enhancements to the attention-based architecture. Lastly, there has been a recent exploration into a probabilistic interpretation of set-to-vector transformations [Zhou et al., 2023, Kim, 2022], further refining the methodology.
@@
The works mentioned above diverge from our approach for several reasons. Firstly, they primarily focus on supervised learning oracles utilizing static datasets to learn neural parameters, whereas our method draws inspiration from max-margin contrastive learning and employs a dynamically updated dataset. Secondly, these existing works lack architectural and/or optimization constraints ensuring the submodularity of the learned set function, in contrast to our work on learning deep submodular functions, where submodularity is explicitly guaranteed. However, for a comprehensive evaluation against existing literature, we include Deep Sets and Set-Transformer as a baseline for comparison. Thirdly, the set sizes used in these works are much smaller (30 in many experiments) compared to ours which go up to 500. Lastly, the learned set function in these works lacks a clear interpretation associated with extremal—sets that (locally) maximize or minimize the set function. In contrast, our Deep Submodular Function (DSF) yields heterogeneous/homogeneous sets upon maximization/minimization. It is noteworthy that the choice of a submodular function as the inductive bias may not be essential for certain tasks, such as point-cloud classification and amortized clustering of mixture models. However, it proves to be a plausible inductive bias in scenarios where the objective is to count unique characters [Zaheer et al., 2017, Zhou et al., 2023, Kim, 2022, Lee et al., 2019], since it can be exactly be expressed as a partition matroid rank function, which lies in the family of deep submodular function.
@@
Beyond neural approaches, one of the previous works [Mariet et al., 2019] has also considered learning Determinantal Point Process, or DPPs [Kulesza et al., 2012], by learning the $\mathcal{O}(n^{2})$ similarity matrix by noise contrastive estimation [Gutmann and Hyvärinen, 2012]. However, their learning procedure is different from ours, and sampling from a DPP is much more computationally expensive than maximizing a submodular function.
Theoretical Work on Learning Submodular Functions
@@
Numerous works have investigated strategies for understanding submodular functions from a theoretical perspective. Goemans et al. [2009] examined the approximation of an oracle for every subset, utilizing a multiplicative approximation approach to approximate the submodular polyhedron of the oracle using Löwner ellipsoids. In contrast, Balcan et al. [2011], Balcan and Harvey [2018] focused on approximating the oracle with high probability on sets sampled from specific distributions (often referred to as Probabilistic Modular Agnostic (PMAC) learning) by transforming the problem into the identification of linear separators. Notably, Balcan and Harvey [2018] demonstrated the impracticality of approximating any arbitrary monotone non-decreasing submodular function to arbitrary closeness under PMAC learning in polynomial queries. Subsequently, Balcan et al. [2016] proposed an algorithm for learning the oracle through pairwise comparisons in polynomial queries. While this work shares theoretical commonalities with ours, it is important to highlight that Balcan et al. [2016] examined set pairs sampled from a static distribution, contrasting with our approach, which involves a dynamically changing distribution of set pairs due to our submodular feedback. Remarkably, across the aforementioned works, the learned functions exhibit a form that can be represented easily using a Deep Submodular Function (DSF), a facet that our work incorporates and explores further.
Neural Estimation of Submodular Functions
@@
Since submodular functions have $2^{|V|}$ degrees of freedom, empirical approaches of learning submodular functions restrict the search space to a parametric subclass such as feature-based functions or DSFs. Early data-driven learning techniques use the max-margin structured prediction framework introduced in Taskar et al. [2005], which requires human annotators to manually construct summaries/sets that should be assigned high values by the learnt submodular function [Lin and Bilmes, 2012, Tschiatschek et al., 2014, Sipos et al., 2012, Gygli et al., 2015, Ghadimi and Beigy, 2020]. These approaches have consistently demonstrated that learned coefficients, as opposed to handcrafted ones, result in submodular functions that can generate much higher quality summaries. However, these approaches depend on human-constructed summaries which can be prohibitively expensive to procure at scale. Furthermore, that the features suitable for the feature-based function are not learned, unlike our work where we jointly learn features and associated parameters for the submodular function. Despite the empirical success of this framework, these approaches all require sets chosen by humans which are difficult to procure at scale.
@@
More recently, De and Chakrabarti [2022] proposed to estimate a submodular function using recursive composition with a learned concave function, arguing that the lack of a “good" concave function is a weakness of DSF (and only used one concave function in their experiments). In our work, we alleviate this by showing that one can learn a good DSF by training with multiple concave functions simultaneously with varying degrees of saturation. It should also be noted that there is no theoretical discussion if the representation capacity of the function class strictly increases with more recursive steps, unlike DSF, where the capacity strictly increases with additional layers. Furthermore, they learn the parameters in a supervised manner, given the annotated high-valued summaries, unlike our work which does so in a self-supervised fashion. Lastly, the learned modular function for DSF has the interpretation of capturing concepts (appendix I) which is not in their case. Remarkably, it is an interesting extension where one combines the technique proposed in De and Chakrabarti [2022] to learn the concave function jointly with our technique for learning the parameters.
@@
Lavania and Bilmes [2019] proposes an unsupervised framework to learn submodular functions, similar to our work. Their approach involves the maximization of a handcrafted objective function, evaluating the quality of weights through measurements of submodular function properties like curvature and sharpness. Notably, they do not engage in the simultaneous joint learning of features alongside parameters and instead they first train an auto-encoder to get the features. To the best of our knowledge, no existing studies have effectively undertaken the joint learning of submodular function parameters and features, a key aspect of our proposed methodology.
@@
Finally, we note the existence of related work using the same nomenclature, DSPN, introduced by Zhang et al. [2019], which stands for Deep Set Prediction Network. The primary objective of Zhang et al. [2019] is to predict a set based on dense features. In contrast, our research focuses on the development of a deep submodular function (DSF) capable of generating sets through optimization.

### v3

@@ 5.2 Application to Experimental Design
Experimental design involves selecting a subset of unlabeled examples from a large pool for labeling to create a training set [80]. In this section, we evaluate the performance of a learnt DSPN based on how well a simple linear model generalizes when trained on a training set that was chosen by a DSPN. We consider an offline setting where we assume the full unlabeled pool is available and an online setting where the unlabeled pool is presented to the model in a non-iid stream. To test the robustness of our framework to real-world challenges, we add duplicates such that the ground set is heavily class-imbalanced (the procedure to generate the class-imbalanced set is shown in Appendix H). Importantly, in these experiments we train with one set $V$ and test on a completely held-out set $V^{\prime}$ with $V\cap V^{\prime}=\emptyset$; that is, the testing DSPN is on an entirely different ground set than the training DSPN.
@@
Offline Setting: The offline setting assumes that the full unlabeled pool is accessible by the summarization procedure. The results of these experiments are presented in Figure 6. Our approach, denoted as DSPN in the figure, applies greedy maximization to the learnt DSPN to generate the summary. We also consider a cheating experiment by applying greedy to the target FL after removing duplicates from the ground set; we refer to this approach as Target FL. Finally, we consider three baseline summarization procedures CRAIG [69], K-centers [89], and random subset selection [75]. The learnt DSPN achieves performance comparable to the target FL in all datasets up to a budget of 100. Beyond that, the difference between DSPN and the Target FL grows but the DSPN continues to generate higher quality than any of the baselines, showing DSPNs generalize to set sizes beyond the training set sizes.
@@
Online Setting: We consider an online setting where the data is presented to the model in a non-iid, class-incremental stream, as done in the continual learning literature [63; 18; 2; 113]. Unlike the facility location function which requires the full ground set for evaluation, the DSPN can be maximized in a streaming manner. Thus, we employ the streaming maximization algorithm proposed by [52] to maximize the DSPN, due to its effectiveness in real-world settings and its lack of additional memory requirements. This approach is referred to as DSPN-online in Figure 6.
@@
We compare our approach against two streaming summarization baselines. The first approach, reservoir sampling [104], is an algorithm that enables random sampling of a fixed number of items from a data stream as if the sample was drawn from the entire population at once. The second approach, VeSSAL [86], seeks to select samples such that the summary maximizes the determinant of the gram matrix constructed from the model embeddings. We also apply (offline) greedy maximization to the target facility location (Target FL-offline) and the learned DSPN (DSPN-offline), which serve as upper bounds to what is attainable by streaming DSPN maximization.
@@
In Figure 6, we present quantitative results where a linear model is trained on a given streaming summary after labels have been obtained. We find that DSPN-online significantly outperforms the other streaming baselines on all datasets. Surprisingly, DSPN-offline and DSPN-online achieve similar performance, demonstrating the flexibility of DSPNs. FL-offline tends to generate the highest quality summaries, but requires access to the full ground set and assumes duplicates are manually removed.
@@ 5.3 Ablations
Table 1: Ablations on Imagenet100. The importance of feedback, learnt features, enforcing submodularity, and GPC (in each column, bold is best, underline is 2nd best). “w/o feedback" indicates a DSPN model trained without feedback, while “w/o target feedback” is when the oracle is only queried. For Set Transformer, streaming maximization fails to produce a summary of the required size (therefore it is omitted). “PC rather than GPC” uses non-graded pairwise comparisons, showing the benefits of grading, consistent with [61]. All results are for a size-100 set.
Table 1, provides Imagenet100 results on sets of size 100 that explore the importance of various components of the DSPN. We report three numbers for each ablation: (1) FL Eval which corresponds to the normalized FL evaluation discussed in § 5.1 (2) Offline Acc which denotes the final accuracy of a linear model in the offline experimental design setting and (3) Online Acc denotes the final accuracy of a linear model in the online experimental design setting.
Removing (FL) Feedback: Active selection is obtained by maximizing the DSPN or Target FL during training (denoted as DSPN feedback and Target feedback respectively). We find that only removing the sets obtained from target feedback has little to no impact on the final performance of the DSPN, suggesting that it is not necessary to optimize the target function. This shows that our framework can be applied to GPC learning even when it is not possible to optimize the target, such as with human oracles. However, we find that completely removing Active selection significantly hurts the final DSPN .
Removing Learnable Pillar: We assess the utility of a learnable pillar, by training a DSPN with a fully frozen pillar. In this case, we attempt to learn the weights of a DSF on random projections of CLIP features. We discover that the DSPN without learnable pillars significantly underperforms its counterpart. We discuss this further in appendix I .
Removing Submodularity: We study the utility of submodular constraints to learn the target. To this end, we compare DSPN against a Deep Set [110] and SetTransformer [55] architecture in the third section of table 1. we discover that the absence of submodularity significantly hampers the performance for all the metrics, since the greedy algorithm has no guarantees on an arbitrary set function.
@@ 6 Conclusion and Future Work
We have proposed DSPNs and methods for training/distilling them. Since the computational cost for querying a DSPN is independent of dataset size, and the training scheme leverages automatically generated $(E,M)$ pairs, our methods learn practical submodular functions scalably. It may ultimately be possible to scale our training framework to massive heterogeneous datasets to develop foundation DSPNs having strong zero-shot summarization capabilities on a wide array of domains. We also propose the peripteral loss which leverages numerically graded relationships between pairs of objects. The graded pairwise preferences are provided from an oracle and are used to learn a submodular function that captures these relationships. However, the scope of the peripteral loss extends far beyond training DSPNs and has many applications such as new contrastive representation learning procedures, and learning a reward model in the context of RLHF with real human oracles.
7 Acknowledgments
We thank Suraj Kothawade, Krishnateja Killamsetty and Rishabh Iyer for several early discussions. This work is supported in part by the CONIX Research Center, one of six centers in JUMP, a Semiconductor Research Corporation (SRC) program sponsored by DARPA and by NSF Grant Nos. IIS-2106937 and IIS-2148367.
@@ Appendix A Other Related Work
A.1 Relationship to contrastive learning
We are inspired by contrastive learning [8], which has pushed the boundaries of deep learning by leveraging the massive information in unlabeled data. Unlike standard supervised learning which teaches models to associate samples to human-supplied labels, contrastive learning encourages models to capture relationships between pairs, groups, or sections of unlabeled samples. It does this by “contrasting” positive pairs (i.e., similar data points) against negative pairs (dissimilar data points). Mathematically, this approach uses loss functions that encourage the model to learn embeddings where positive pairs become “close” (according to some distance measure), while negative pairs become “far.” A common setting is the contrastive loss, such as the triplet loss or the Siamese network loss. For instance, in the triplet loss framework, the model is trained on triplets of data points consisting of an anchor $x$, a positive example $x^{+}$ (similar to the anchor), and a negative example $x^{-}$ (dissimilar from the anchor). Optimizing a model with such a loss function encourages a representation that respects these positive/negative pairwise preferences.
On the other hand, such contrastive approaches are analogous to binary-only pairwise comparison elicitations. That is, existing contrastive losses neither utilize information regarding how positive the pair $x,x^{+}$ is nor how negative the pair $x,x^{-}$ is. A more nuanced graded comparison, if available, would be an improvement. In fact, the benefit of graded comparisons is analogous to the benefits of using soft over hard labels in supervised training (say from model distillation) which has benefits in uncertainty representation, regularization, noise mitigation, adjustable smoothness via softmax temperature, and more efficient knowledge transfer. Another limitation of previous contrastive learning is that it operates only with positive or negative groups of size two (pairs) rather than larger groups. While it is often the case that the anchor $x$ is contrasted with many negative examples, it is always done so pairwise, considering each $x^{-}$ with $x$ together in a sum of pairwise terms. Also, while [100, 42] is the only work we are aware of that considers more than two positive examples at a time, each of these positive samples considered are only paired up with a negative sample, leading to $\binom{k}{2}$ pairings with $k$ positive examples. Sums of pairwise terms do not allow for the consideration of higher order relationships, e.g., a given pair might be measured differently in the context of a third sample. In addition to addressing the two aforementioned problems, DSPNs and our GPC-style loss addresses this issue as well, although our learning goal in this paper is learning submodularity not in representation learning (the latter of which we leave to future work).
A.2 Learning Non-submodular Set functions
@@
Recently, there has been a growing focus on leveraging data-driven neural methods for estimating set functions, with practical applications spanning anomaly detection, amortized clustering of mixture models, and point cloud classification [81]. Common to these applications is the fundamental approach of embedding each element within a set into dense embeddings, followed by permutation-invariant aggregation (set-to-vector transformation), and subsequent integration into a deep neural network [110, 38, 83]. The significance of incorporating interaction among set elements has been emphasized by [55], who argue that conventional aggregations, such as average-pooling, can be interpreted as a specialized form of attention. Consequently, they propose an attention-based architecture known as the set-transformer. Building upon this, [67] have made computational enhancements to the attention-based architecture. Lastly, there has been a recent exploration into a probabilistic interpretation of set-to-vector transformations [113, 46], further refining the methodology.
@@
The works mentioned above diverge from our approach for several reasons. Firstly, they primarily focus on supervised learning oracles utilizing static datasets to learn neural parameters, whereas our method draws inspiration from max-margin contrastive learning and employs a dynamically updated dataset. Secondly, these existing works lack architectural and/or optimization constraints ensuring the submodularity of the learned set function, in contrast to our work on learning deep submodular functions, where submodularity is explicitly guaranteed. However, for a comprehensive evaluation against existing literature, we include Deep Sets and Set-Transformer as a baseline for comparison. Thirdly, the set sizes used in these works are much smaller (30 in many experiments) compared to ours which go up to 500. Lastly, the learned set function in these works lacks a clear interpretation associated with extremal—sets that (locally) maximize or minimize the set function. In contrast, our Deep Submodular Function (DSF) yields heterogeneous/homogeneous sets upon maximization/minimization. It is noteworthy that the choice of a submodular function as the inductive bias may not be essential for certain tasks, such as point-cloud classification and amortized clustering of mixture models. However, it proves to be a plausible inductive bias in scenarios where the objective is to count unique characters [110, 113, 46, 55], since it can be exactly be expressed as a partition matroid rank function, which lies in the family of deep submodular function.
@@
Beyond neural approaches, one of the previous works [66] has also considered learning Determinantal Point Process, or DPPs [50], by learning the $\mathcal{O}(n^{2})$ similarity matrix by noise contrastive estimation [35]. However, their learning procedure is different from ours, and sampling from a DPP is much more computationally expensive than maximizing a submodular function.
A.3 Theoretical Work on Learning Submodular Functions
@@
Numerous works have investigated strategies for understanding submodular functions from a theoretical perspective. [31] examined the approximation of an oracle for every subset, utilizing a multiplicative approximation approach to approximate the submodular polyhedron of the oracle using Löwner ellipsoids. In contrast, [5, 6] focused on approximating the oracle with high probability on sets sampled from specific distributions (often referred to as Probably Mostly Approximately Correct (PMAC) learning) by transforming the problem into the identification of linear separators. Notably, [6] demonstrated the impracticality of approximating any arbitrary monotone non-decreasing submodular function to arbitrary closeness under PMAC learning in polynomial queries. Subsequently, [7] proposed an algorithm for learning the oracle through pairwise comparisons in polynomial queries. While this work shares theoretical commonalities with ours, it is important to highlight that [7] examined set pairs sampled from a static distribution, contrasting with our approach, which involves a dynamically changing distribution of set pairs due to our submodular feedback. Remarkably, across the aforementioned works, the learned functions exhibit a form that can be represented easily using a Deep Submodular Function (DSF), a facet that our work incorporates and explores further.
A.4 Neural Estimation of Submodular Functions
@@
Since submodular functions have $2^{|V|}$ degrees of freedom, empirical approaches of learning submodular functions restrict the search space to a parametric subclass such as feature-based functions or DSFs. Early data-driven learning techniques use the max-margin structured prediction framework introduced in [98], which requires human annotators to manually construct summaries/sets that should be assigned high values by the learnt submodular function [60, 102, 93, 36, 29]. These approaches have consistently demonstrated that learned coefficients, as opposed to handcrafted ones, result in submodular functions that can generate much higher quality summaries. However, these approaches depend on human-constructed summaries which can be prohibitively expensive to procure at scale. Furthermore, that the features suitable for the feature-based function are not learned, unlike our work where we jointly learn features and associated parameters for the submodular function. Despite the empirical success of this framework, these approaches all require sets chosen by humans which are difficult to procure at scale.
@@
More recently, [22] proposed to estimate a submodular function using recursive composition with a learned concave function, arguing that the lack of a “good" concave function is a weakness of DSF (and only used one concave function in their experiments). In our work, we alleviate this by showing that one can learn a good DSF by training with multiple concave functions simultaneously with varying degrees of saturation. It should also be noted that there is no theoretical discussion if the representation capacity of the function class strictly increases with more recursive steps, unlike DSF, where the capacity strictly increases with additional layers. Furthermore, they learn the parameters in a supervised manner, given the annotated high-valued summaries, unlike our work which does so in a self-supervised fashion. Lastly, the learned modular function for DSF has the interpretation of capturing concepts (appendix I) which is not in their case. Remarkably, it is an interesting extension where one combines the technique proposed in [22] to learn the concave function jointly with our technique for learning the parameters.
@@
[51] proposes an unsupervised framework to learn submodular functions, similar to our work. Their approach involves the maximization of a handcrafted objective function, evaluating the quality of weights through measurements of submodular function properties like curvature and sharpness. Notably, they do not engage in the simultaneous joint learning of features alongside parameters and instead they first train an auto-encoder to get the features. To the best of our knowledge, no existing studies have effectively undertaken the joint learning of submodular function parameters and features, a key aspect of our proposed methodology.
@@
Finally, we note the existence of related work using the same nomenclature, DSPN, introduced by [112], which stands for Deep Set Prediction Network. The primary objective of [112] is to predict a set based on dense features. In contrast, our research focuses on the development of a deep submodular function (DSF) capable of generating sets through optimization.